OpenStudy (anonymous):

im having trouble integrating the double iterated integral R= [0,3] X [0,rad(9-y^2)] and the function inside the integral is "y dy dx"

OpenStudy (anonymous):

\[\int\limits_{0}^{3}\int\limits_{0}^{\sqrt{9-7^{2}}} y dy dx\] thats the integral

OpenStudy (anonymous):

something's not right with either your limits or your order of integration...

OpenStudy (anonymous):

sorry your right.... its y dx dy

OpenStudy (anonymous):

if you're integrating in 'dy' .. your limits will be functions of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

that works...

OpenStudy (anonymous):

integral of ydx is yx +C

OpenStudy (anonymous):

ive tried u sub...and integration by parts. but i just cant get the answer....

OpenStudy (anonymous):

still something weird about it though... shouldn't be written like that...

OpenStudy (anonymous):

the problem just says.... evaluate the iterated integral

OpenStudy (anonymous):

k

OpenStudy (anonymous):

they're writing it weird to try and throw you off I guess...

OpenStudy (anonymous):

anyway... plug in the limits on 'x' ... ( 0 ... srt(9-y^2) ) and then evaluate with respect to dy

OpenStudy (anonymous):

but then what? dont we need to simplifly

OpenStudy (anonymous):

so (y*sqrt(9-y^2) + C) dy

OpenStudy (anonymous):

you can use u sub.s ...

OpenStudy (anonymous):

9-y^2 = u -2y dy = du y dy = -du/2

OpenStudy (anonymous):

yes but then we need to plug in 3 into \[\sqrt{9-3^{2}}\] wich equals 0 in the rad

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

ends up canceling everything out.... and the answer is supposed to be 3

OpenStudy (anonymous):

...

OpenStudy (anonymous):

lower limit?

OpenStudy (anonymous):

is 0

OpenStudy (anonymous):

omg....

OpenStudy (anonymous):

ive done this problem like no lie 10 times and i never pluged in the lower limit = 0 giving me the right answer 9

OpenStudy (anonymous):

thank you so much for letting me realize this. how to i give a medal??

OpenStudy (anonymous):

click 'best response'